In a simple firing experiment, 1.03g of an explosive material, which can produce a gas product up to 9.2 dm3 per gram if detonated, was packed in a handgun ammunition cartridge and then triggered to detonate to propel a projectile weighed 2 g out of the barrel. The bullet velocity for this firing is 284 m/s and the factor η for the inevitable inefficiency of the energy transfer from the explosion to the projectile is 0.13. Calculate the power index for this explosive material by comparing to the picric acid, which can evolve 3150 J/g of heat and produces 0.831 dm3 of gas for every gram of detonation.


1
Expert's answer
2020-12-01T14:26:21-0500

Energy of projectile = 0.002(2g = 0.002kg) * 2842 /2 = 80.656 J If inefficiency of the energy transfer - 0.13, than energy of explosion was 80.656/0.13 = 620.43 J, what give us 602.43/1.03 = 602.36 J/g; Power index for this explosive material by comparing to the picric acid = Q*V/Q(picric)*V(picric) = 620.43* 9.2/3150 * 0.831 = 1.506